# VLMs have Tunnel Vision: Evaluating Nonlocal Visual Reasoning in Leading VLMs

### v1

@@ 4.2 Visual Scavenger Hunt
The results for Visual Scavenger Hunt are shown in Figure 4. Gemini 2.5 Pro, o4-mini, and o3 are able to perform well above random accuracy of  9.00%. o4-mini shows a clear trend downward as chain length increases. For all models except o4-mini and Gemini 2.5 Pro performance is static with respect to the length of the chain. Claude 3.7 Sonnet, Phi-4-Multimodal, o3, and Gemma 27b perform a few percentage points above this baseline. The rest of the models score at about the accuracy of random guessing.
Figure 4: Accuracy on the Visual Scavenger Hunt task. Only o3, o4-mini, and Gemini 2.5 Pro significantly outperform random chance.
We manually observe the first 20 responses of all models. The weaker models exhibit guessing behavior. They hallucinate paths even for a chain length of 2; an example of this type of response can be found in Figure 5. Even with correct final colors, justifications often cite non-existent shapes or paths. o4-mini often claims it cannot read the text and refuses to answer the question. We debunk this claim with an ad-hoc study; in isolation, o4-mini was able to extract text out of a randomly selected tile on 20 random examples from our dataset. However, it cannot always accurately extract the entire grid, as shown in Figure 5. It successfully extracts every label, but hallucinates when it transcribes the actual shapes and colors. In our empirical analysis, all of Gemini 2.5 Pro’s mistakes occurred mid-search and were because it confuses a nearby slot that has a similar color, label, or shape.
@@
Figure 5: Example responses for Visual Scavenger Hunt from our qualitative analysis. Most models can locate the first shape but have trouble extending the chain from there. o4-mini and Gemini 2.5 Pro are both high performing but use different strategies.
@@
Gemini 2.5 Pro and o4-mini lack the ability to self-correct in this trial. Although examples with longer chains have more opportunity for error, they also provide additional signal that a mistake has been made. The other shapes in the image are labeled randomly, meaning about 30% of the time they lead to a shape that does not exist on the grid. For a human performing the task, this would be a strong signal of a prior mistake, but we do not observe this process across any model.
@@ 4.3 Circuit Connections
Figure 6: Results on the Circuit Connections task. Performance rises across the board from single color to unique, suggesting that heuristics that do not involve tracing the line play a large role in model success.
The human solution to this task relies on tracing each wire from its source to its target. Models could instead use a simple color-matching shortcut: identify the port’s wire color, then pick at random one of the components wired in that color. This strategy alone raises accuracy well above the 14.3% chance baseline and requires no true contour tracking.
Table 1 presents the log odds coefficients for two key characteristics: the Euclidean distance of the connecting wire and the number of times this wire crosses other wires. To understand the model strategy for solving this problem, we measured how each additional pixel of Euclidean distance from the source port to the destination component and each extra crossing altered the log-odds of a correct answer. We hypothesize that for a human-like tracing process, each pixel of extra distance should lower the likelihood of a correct response, and each additional crossing should impose its own penalty. However, this effect may be quite subtle, especially in the low performance regime. We thus look at the difference in log odds across trials. Across trials we predicted a gradient for a model performing Smooth Visual Search: the Single Color variant would show the steepest decline in performance per pixel and per crossing, and the Unique Colors variant would show the smallest decline.
@@
Gemini 2.5 Pro -0.0084 0.00546 -0.5349 0.0244 -0.0086 $6.45\times 10^{-5}$0.1656 0.204 -0.0055 0.0333 -0.2761 0.151
o4-mini -0.0186 $7.81\times 10^{-5}$0.7154 0.0312 -0.0102 $1.1\times 10^{-5}$0.3925 0.0142 -0.0033 0.189 -0.0519 0.782
Claude 3.7 Sonnet -0.0106 0.00328 -0.9305 0.0102 -0.0092 $4.62\times 10^{-5}$0.2021 0.153 -0.0097 $6.94\times 10^{-4}$0.1130 0.561
InternVL3 (14B) -0.0053 0.114 -0.0587 0.798 -0.0062 $3.63\times 10^{-3}$ +0.0727 0.567 -0.0121 $1.95\times 10^{-4}$0.3154 0.191
Phi-4 Multimodal (14B) +0.0008 0.791 -0.0004 0.999 -0.0070 $8.2\times 10^{-3}$0.0935 0.595 -0.0015 0.594 +0.0193 0.931
Qwen2.5VL (32B) -0.0053 0.114 -0.0587 0.798 -0.0137 $2.63\times 10^{-6}$0.1847 0.291 -0.0080 0.0129 -0.8284 0.0209
Qwen2.5VL (7B) -0.0065 0.0692 -0.2718 0.308 -0.0042 0.0839 +0.0236 0.878 -0.0083 $8.4\times 10^{-3}$0.2162 0.386
Mistral Small 3.1(24B) -0.0160 0.00103 -0.2182 0.449 -0.0108 $4.53\times 10^{-5}$0.2331 0.188 -0.0081 $9.18\times 10^{-3}$0.2546 0.309
Gemma 3 (27B) -0.0133 0.00321 -0.5133 0.136 -0.0165 $5.01\times 10^{-7}$0.1152 0.509 -0.0121 $8.13\times 10^{-4}$0.4617 0.139
@@
Llama-3.2 Vision (11B) -0.0020 0.521 -0.4820 0.0868 -0.0141 $6.07\times 10^{-5}$0.2813 0.252 -0.0079 0.0303 -0.1177 0.683
o3 -0.0102 $1.25\times 10^{-5}$0.3005 0.049
Heuristic Strategies
These results imply that high-performing VLMs have increased their visual resolution but they are incapable of following a contour consistently. The highest performing models– Gemini 2.5, o4-mini, Claude 3.7, and to a lesser extent, InternVL– lose their performance edge on the Single Color trial. Their relatively high performance when every wire color is unique indicates they are at least adept at using general context to figure out where the wire goes.
Euclidean distance from the port to the target component is the strongest predictor of model accuracy across all three trial types. In both the Standard and Unique Colors variants, nearly every model shows a significant negative distance coefficient. Only our top performers—Gemini 2.5 Pro, o4-mini and Claude 3.7 Sonnet—consistently incur a penalty for extra crossings. Lower-performing models exhibit little to no evidence for a strong crossings effect and sometimes show no distance penalty when color cues alone suffice.
Our data suggest that most models cannot trace lines and is inconclusive for the top performers. Although nearly every model outperforms random chance, the log-odds analysis shows non-significant crossings effects for the majority. These patterns, along with low overall accuracy, point to a reliance on coarse spatial and color heuristics. In contrast, Gemini 2.5 Pro, o4-mini, and Claude 3.7 Sonnet roughly follow our predicted gradient across variants. However, their absolute accuracy is very poor. As a result, we cannot determine whether they possess an inaccurate contour-tracing mechanism or simply apply more sophisticated heuristics.
@@ 5 Conclusion
We introduce a suite of targeted tasks to distill nonlocal visual reasoning into three components: comparative perception, saccadic search, and smooth visual search. Most VLMs tested fail each benchmark, even those designed to deal with structured data. The flagship closed-source VLMs achieved the highest accuracy but perform poorly relative to humans, which suggests they not only lack basic perception skills.
These findings strongly argue for a shift in focus towards models that explicitly support structured and systematic visual reasoning that is independent of natural language. By exposing these failure modes, we aim to steer the development of VLMs towards architectures that not only describe visual scenes but can genuinely think over the pixels they process.
6 Acknowledgments
This work was partially supported by the National Science Foundation.
@@ Appendix A Task Details: Examples and Prompts
For clarity, we provide visual examples of the task pulled directly out of our dataset..
@@ A.1 Object Re-Identification
Figure 7: Examples of the Object Re-Identification (Standard Variant) task. Prompt: "The first image shows an object made of connected geometric shapes, which together form an object. Does this SAME object appear in the second image? For example, if a component shape were to be rotated or translated separately from the entire composite-object, it would be a different object. Respond with {yes} or {no} (inside the curly brackets). There may be extra shapes in Image 2 that are not part of the original object; as long as the object from Image 1 is present, the answer is yes even if there are other shapes present." [Two examples with answers would precede this.]
@@
Figure 8: Examples of the Object Re-Identification (Unconnected Variant) task. Prompt: "The first image shows an object made of geometric shapes, which together form an object. Does this SAME object appear in the second image up to a rigid translation, rotation, and scale of the ENTIRE object as a whole? Respond with {yes} or {no}. There may be extra shapes in Image 2 that are not part of the original object; as long as the object from Image 1 is present, the answer is yes even if there are other shapes present." [Two examples with answers would precede this.]
@@
Figure 9: Examples of the Object Re-Identification (Pixel-Perfect Variant) task. Prompt: "The first image shows an object made of geometric shapes, which together form an object. Does this SAME object appear in the second image? For example, if a component shape were to be rotated or translated, it would be a different object. Respond with {yes} or {no} (inside the curly brackets). There may be extra shapes in Image 2 that are not part of the original object; as long as the object from Image 1 is present, the answer is yes even if there are other shapes present." [Two examples with answers would precede this.]
@@ A.2 Visual Scavenger Hunt
Figure 10: Example of the Visual Scavenger Hunt task. For a chain length of 4, the prompt would be: “Starting at the blue square, follow the labels for 4 steps. (For instance, in a different example of 4 steps, you might start at a blue triangle, then go to a red square, then a blue circle, then a magenta triangle, then a green circle. The answer would be green.) After those steps, what color are you on? Answer with the color in curly braces, e.g. {red}." [A single example tracing the path on a different image would preced this.]
OCR Capability Confirmation Prompts (Section 3.2):
To confirm that the text was readable, we tested Qwen 2.5VL (7B), o4-mini-high, and Llama 3.2 Vision (11B) on 20 examples from our variant of chain length 3. We used the following prompt: “What does the text in the top-left shape say? " Each model scored perfectly, as seen in Table 2.
Table 2: OCR Capability Test Results
Model Name Score
Qwen 2.5VL (7B) 20/20
Llama 3.2 Vision (11B) 20/20
o4-mini-high 20/20
@@ A.3 Circuit Connections
The Circuit Connections task (Section 3.3) assesses smooth visual search, requiring the model to trace a wire from a specified port on a central breadboard to its connected component. Examples for its three variants, based on wire coloring, are shown in Figure 11. The subfigures show the Standard variant (Figure 11a), Single Color variant (Figure 11b), and Unique Colors variant (Figure 11c).
@@
Figure 11: Examples of the Circuit Connections task variants. The prompt: “Which component does the wire from port 7 on the breadboard, which is the gray rectangle with numbered ports, connect to? A wire is a series of connected, same colored lines that go from the center of a port, represented on the screen as a white circle, to another port. Each wire only connects two ports, one at either end. A wire will NEVER turn at the same spot that it intersects another wire, and wires do not change colors. Answer with the component label in curly braces, e.g {C0}." [Two examples would precede this.]
@@ B.1 Evaluation Parameters
We evaluate all models on highest-reasoning setting and a temperature of $0.01$ where applicable. The answers are normalized and extracted from model responses in plain-text space. Each main trial— the Standard variant of Object Re-Identification and Circuit Connections and Visual Scavenger Hunt— consists of 200 examples and all others are composed of 125 examples each. To control for the possibility that the models do not understand the instructions, we evaluate in a few-shot setting across all experiments and variants. In Object Re-Identification, we provide one positive and one negative example. In Visual Scavenger Hunt we provide a single example which traces the path. In Circuit Connections we provide two examples with the answers. We provide the same example(s) across each trial and for models.
@@
B.2 Computational Resources
Experiments were conducted using a combination of local high-performance computing resources and commercially available model APIs. For the open-source models, we evaluated locally on a system equipped with three NVIDIA A6000 GPUs (48GB VRAM each) and dual Intel(R) Xeon(R) Gold 5220R CPUs @ 2.20GHz. For other models, we accessed them via API. We evaluated Claude 3.7 Sonnet and Gemini 2.5 Pro via OpenRouter at model codes anthropic/claude-3.7-sonnet:thinking and google/gemini-2.5-pro-preview. These models were evaluated between May 1st, 2025 and May 15th, 2025. The OpenAI models were evaluated via the OpenAI API at model codes o4-mini-2025-04-16 and o3-2025-04-16. We gathered model responses between April 28, 2025, and May 14, 2025.

### v2

@@ 4.2 Visual Scavenger Hunt
The results for Visual Scavenger Hunt are shown in Figure 4. Gemini 2.5 Pro, o4-mini, and GPT-5 perform well above the random-guess baseline of  9%. o4-mini shows a clear trend downward as chain length increases.For all models except GPT-5, o4-mini, and Gemini 2.5 Pro, performance is mostly static with respect to the length of the chain. Claude 3.7 Sonnet, Phi-4-Multimodal, o3, and Gemma 3 27b perform a few percentage points above this baseline. The rest of the models score at about the accuracy of random guessing.
Figure 4: Accuracy on the Visual Scavenger Hunt task. Only GPT-5, o4-mini, and Gemini 2.5 Pro significantly outperform random chance.
Failure Modes
We manually observe the first 20 responses of all models. The weaker models exhibit guessing behavior. They hallucinate paths even for a chain length of 2; an example of this type of response can be found in Figure 5. Even with correct final colors, justifications often cite non-existent shapes or paths. o4-mini often claims it cannot read the text and refuses to answer the question. We debunk this claim with an ad-hoc study; in isolation, o4-mini was able to extract text out of a randomly selected tile on 20 random examples from our dataset. However, it cannot always accurately extract the entire grid, as shown in Figure 5. It successfully extracts every label, but hallucinates when it transcribes the actual shapes and colors. In our empirical analysis, all of Gemini 2.5 Pro’s mistakes occurred mid-search and seemed to be due to confusion with a nearby slot that had a similar color, label, or shape.
To determine if failures on this task stem from an inability to locate the shapes and text, or from a breakdown in chaining these steps through saccadic search, we performed a follow-up experiment. We decomposed the chain length 3 task into three sequential, single-step queries, where the model’s answer for one step was used as the starting point for the next. The results are shown in Table 1. In addition to overall accuracy on the three-step sequence, we report the Final-Step Error Rate. This metric is calculated only on trials where the model correctly navigated the first two steps, thereby isolating the error rate of the final jump. We use this conditional metric because grading intermediate steps requires verifying both shape and color, whereas the final step, like our main task, only requires identifying the correct color.
Table 1: Performance on the decomposed Visual Scavenger Hunt task over 75 trials. Accuracy reflects the success rate over the full three-step interactive sequence, while the Final-Step Error is conditioned on the first two steps being correct.
Model Accuracy (%) Final-Step Error (%)
o4-mini 90.67 6
Gemini 2.5 88.00 8
Llama 3.2 Vision (11B) 24.00 85
Qwen 2.5 VL (32B) 16.00 86
Qwen 2.5 VL (7B) 14.67 85
These results show that top-performing models are highly capable of executing the atomic unit of the task when guided, whereas the other models cannot perceive well enough to perform the task. However, their high single-step accuracy contrasts with their low accuracy over a multi-step search. Simple error multiplication suggests o4-mini should achieve  83% accuracy ($0.94^{3}$}), far higher than the 44.5% observed in Figure 4. This discrepancy indicates that the primary bottleneck for the strong models is not perception, but the inability to perform multiple steps autonomously. Conversely, the other models’ poor performance suggests they lack the basic perceptive abilities required to perform a single step correctly.
@@
Figure 5: Example responses for Visual Scavenger Hunt from our qualitative analysis. Most models can locate the first shape but have trouble extending the chain from there. o4-mini and Gemini 2.5 Pro are both high-performing but use different strategies.
@@
Gemini 2.5 Pro and o4-mini lack the ability to self-correct in this trial. Although examples with longer chains have more opportunities for error, they also provide additional signal that a mistake has been made. The other shapes in the image are labeled randomly, meaning about 30% of the time they lead to a shape that does not exist on the grid. For a human performing the task, this would be a strong signal of a prior mistake, but we do not observe this process across any model.
@@ 4.3 Circuit Connections
Figure 6: Results on the Circuit Connections task. Performance rises drastically across the board from Single Color to Unique Colors, suggesting that heuristics that do not involve tracing the line play a large role in model success.
Table 2 presents the log odds coefficients for two key characteristics: the Euclidean distance of the connecting wire and the number of times this wire crosses other wires. The log-odds analysis allows us to interpret by what factor the odds of a binary outcome (e.g., success vs. failure) change for each unit increase in a predictor variable, like the distance between the source and target port or the number of wire-crossings.
@@
GPT-5 -0.0106 0.00131 -1.0008 0.00238 -0.0063 0.00225 +0.0028 0.982 -0.0026 0.299 -0.3477 0.0742
Gemini 2.5 Pro -0.0084 0.00546 -0.5349 0.0244 -0.0086 $6.45\times 10^{-5}$ -0.1656 0.204 -0.0055 0.0333 -0.2761 0.151
o4-mini -0.0186 $7.81\times 10^{-5}$ -0.7154 0.0312 -0.0102 $1.1\times 10^{-5}$ -0.3925 0.0142 -0.0033 0.189 -0.0519 0.782
Claude Sonnet 3.7 -0.0106 0.00328 -0.9305 0.0102 -0.0092 $4.62\times 10^{-5}$ -0.2021 0.153 -0.0097 $6.94\times 10^{-4}$ -0.1130 0.561
Claude Sonnet 4 -0.0134 $6.94\times 10^{-4}$ -0.5653 0.0561 -0.0123 $7.65\times 10^{-7}$ -0.0420 0.749 -0.0097 0.00106 -0.7260 $0.00853$
InternVL3 (14B) -0.0053 0.114 -0.0587 0.798 -0.0062 $0.00363$ +0.0727 0.567 -0.0121 $1.95\times 10^{-4}$ -0.3154 0.191
Phi-4 Multimodal (14B) $8\times 10^{-4}$ 0.791 $4\times 10^{-6}$ 0.999 -0.0070 $0.0082$ -0.0935 0.595 -0.0015 0.594 +0.0193 0.931
Qwen2.5VL (32B) -0.0053 0.114 -0.0587 0.798 -0.0137 $2.63\times 10^{-6}$ -0.1847 0.291 -0.0080 0.0129 -0.8284 0.0209
Qwen2.5VL (7B) -0.0065 0.0692 -0.2718 0.308 -0.0042 0.0839 +0.0236 0.878 -0.0083 $0.0084$ -0.2162 0.386
Mistral Small 3.1 (24B) -0.0160 0.00103 -0.2182 0.449 -0.0108 $4.53\times 10^{-5}$ -0.2331 0.188 -0.0081 $0.00918$ -0.2546 0.309
Gemma 3 (27B) -0.0133 0.00321 -0.5133 0.136 -0.0165 $5.01\times 10^{-7}$ -0.1152 0.509 -0.0121 $8.13\times 10^{-4}$ -0.4617 0.139
@@
Llama-3.2 Vision (11B) -0.0020 0.521 -0.4820 0.0868 -0.0141 $6.07\times 10^{-5}$ -0.2813 0.252 -0.0079 0.0303 -0.1177 0.683
The statistics in Table 2 also suggest that these models are not performing human-like tracing. The strongest predictor of model performance across all trials is distance, with high confidence ($p\leq 0.05$) across almost all models on the Standard and Unique Colors trials. The crossings effect is strongest in the highest-performing models.
Failure Modes
Across all Circuit Connections trials, the models’ success rates and our log odds analysis indicate that no model performs contour tracing. They cluster into two groups— one which does slightly better than random on Unique Colors and Standard, and the three models which significantly beat random accuracy.
Most models we tested show a complete inability to perform line tracing. These models perform at or below random chance (14.29%) across all three variants of the Circuit Connections task (Standard, Single Color, Unique Colors), as depicted in Figure 6.
This widespread low accuracy suggests these models cannot trace lines, and our log-odds analysis in Table 2 supports this assertion. Many models in this group show a significant correlation between Euclidean Distance and success in the Standard or Unique Colors trials when they beat random accuracy. However, the statistical significance of this effect disappears in the Single Color variant (e.g., Molmo (7B) p=0.343) The wire-crossings effect for this group remains largely insignificant across all variants. This general lack of sensitivity to crossings is evidence they do not perform human-like line tracing— intuitively, the task of line-tracing is hardest when two lines interfere. We hypothesize they are able to exceed random accuracy by using color-cues and proximity heuristics.
Top performing models, such as Gemini 2.5 Pro, Claude 3.7 Sonnet, and O4-mini, form a second group that performs better but still demonstrates significant limitations in tracing. Their accuracy drops significantly when color cues are removed or made uniform. For example, Gemini 2.5 Pro’s accuracy falls from 48% on Unique Colors to 27% on Single Color. While tracing is expected to be harder when wires of the same color interfere, this low performance suggests either a lack of, or severely limited, tracing ability. The Euclidean distance of the connecting wire remains a significant negative predictor for these models across all trials. However, what distinguishes them from the first group is that these models do exhibit a statistically significant negative crossing effect in the challenging Single Color trial, where wires often cross other wires of the same color (e.g Claude 3.7 Sonnet shows a log-odds coefficient of -0.9305 (p=0.0102) for the crossings effect in this variant.) This behavior is consistent with a tracing algorithm, where crossings represent visually ambiguous points. However, because these models also do poorly in Unique Colors compared to humans, we suspect these models are possibly performing choppy, saccadic movements to find successive points on the wire rather than a smooth trace. This strategy aligns with the statistics: in the Single Color variant, models get caught in wire crossings and struggle to accurately jump to the next segment, impacting performance.
@@ 5 Conclusion
We introduce a suite of targeted tasks to distill nonlocal visual reasoning into three components: comparative perception, saccadic search, and smooth visual search. Most VLMs tested fail each benchmark, even those designed to deal with structured data. Closed-source VLMs achieved the highest accuracy but performed wildly differently under slight permutations of the underlying task. The task with the lowest performance ceiling was Circuit Connections, which requires smooth visual search. This suggests a principal limitation of these models is their difficulty with tasks that resist natural language reformulation.
This suggests that despite gains on other visual benchmarks, VLMs lack a reliable framework to analyze images. Without these capabilities, the vision of VLMs will remain fundamentally less robust than human vision.
Limitations.
While our synthetic setting isolates primitives, it does not capture the complete gamut of natural images, which may only require visual skims to parse correctly. Additionally, we evaluate on only 200 or 125 examples per variant for cost-efficiency reasons.
@@ Appendix A Task Details: Examples and Prompts
This section provides visual examples for each task category, its variants, and the prompts used.
@@ A.1 Object Re-Identification
Figure 7: Examples of the Object Re-Identification (Standard Variant) task. Prompt: “The first image shows an object made of connected geometric shapes, which together form an object. Does this SAME object appear in the second image? For example, if a component shape were to be rotated or translated separately from the entire composite-object, it would be a different object. Respond with {yes} or {no} (inside the curly brackets). There may be extra shapes in Image 2 that are not part of the original object; as long as the object from Image 1 is present, the answer is yes even if there are other shapes present.” [Two examples with answers would precede this.]
@@
Figure 8: Examples of the Object Re-Identification (Unconnected Variant) task. Prompt: “The first image shows an object made of geometric shapes, which together form an object. Does this SAME object appear in the second image up to a rigid translation, rotation, and scale of the ENTIRE object as a whole? Respond with {yes} or {no}. There may be extra shapes in Image 2 that are not part of the original object; as long as the object from Image 1 is present, the answer is yes even if there are other shapes present.” [Two examples with answers would precede this.]
@@
Figure 9: Examples of the Object Re-Identification (Pixel-Perfect Variant) task. Prompt:“The first image shows an object made of geometric shapes, which together form an object. Does this SAME object appear in the second image? For example, if a component shape were to be rotated or translated, it would be a different object. Respond with {yes} or {no} (inside the curly brackets). There may be extra shapes in Image 2 that are not part of the original object; as long as the object from Image 1 is present, the answer is yes even if there are other shapes present.” [Two examples with answers would precede this.]
@@ A.2 Visual Scavenger Hunt
Figure 10: Example of the Visual Scavenger Hunt task. For a chain length of 4, the prompt would be: “Starting at the blue triangle, follow the labels for 4 steps. (For instance, in a different example of 4 steps, you might start at a blue triangle, then go to a red square, then a blue circle, then a magenta triangle, then a green circle. The answer would be green.) After those steps, what color are you on? Answer with the color in curly braces, e.g. {red}.” [A single example tracing a path (accompanied by an image) would precede this, e.g: “We start at the blue square, and then go to the purple triangle, and then go to the yellow square, and then go to the red square, and then end at the gray square. {gray}” ]
@@ A.3 Circuit Connections
The Circuit Connections task (Section 3.3) assesses smooth visual search, requiring the model to trace a wire from a specified port on a central breadboard to its connected component. Examples for its three variants, based on wire coloring, are shown in Figure 11: the Standard Variant (Figure 11(a)), where multiple wires can share colors; the Single Color Variant (Figure 11(b)), where all wires are the same color; and the Unique Colors Variant (Figure 11(c)), where each wire has a distinct color.
@@
Figure 11: Examples of the Circuit Connections task variants. The prompt: “Which component does the wire from port 7 on the breadboard, which is the gray rectangle with numbered ports, connect to? A wire is a series of connected, same colored lines that go from the center of a port, represented on the screen as a white circle, to another port. Each wire only connects two ports, one at either end. A wire will NEVER turn at the same spot that it intersects another wire, and wires do not change colors. Answer with the component label in curly braces, e.g {C0}.” [Two examples would precede this.]
@@ B.1 Evaluation Parameters
Secondary metrics for o3 were captured incorrectly due to a bug in the evaluation code. They are thus omitted from Figure 3 and Table 2.
B.1.1 Computational Resources
Experiments were conducted using a combination of local high-performance computing resources and commercially available model APIs. For the open-source models, we evaluated locally on a system equipped with three NVIDIA A6000 GPUs (48GB VRAM each) and dual Intel(R) Xeon(R) Gold 5220R CPUs @ 2.20GHz. For other models, we accessed them via API. We evaluated Claude 3.7 Sonnet, Gemini 2.5 Pro, and Claude Sonnet 4 via OpenRouter at model codes anthropic/claude-3.7-sonnet:thinking, google/gemini-2.5-pro-preview, and anthropic/claude-sonnet-4, respectively. These OpenRouter models were evaluated between May 1, 2025 and May 15, 2025 (for Claude 3.7 Sonnet and Gemini 2.5 Pro) and between September 17, 2025 and September 20, 2025 (for Claude Sonnet 4). The OpenAI models were evaluated via the OpenAI API at model codes o4-mini-2025-04-16, o3-2025-04-16, and gpt-5-2025-08-07. We gathered model responses for these evaluations between April 28, 2025 and May 14, 2025 for o4-mini-2025-04-16, between July 24, 2025 and July 27, 2025 for o3-2025-04-16, and between September 17, 2025 and September 20, 2025 for gpt-5-2025-08-07.